OpenStudy (cxybvby):

1/2x + 3y = 1 x + 6y = 2 A. Infinity solutions B. No solutions C. One solution

satellite73 (satellite73):

multiply the first equation by 2 (all the way across) what do you get?

Directrix (directrix):

1/2x + 3y = 1 2 * (1/2) x + 2 * 3y = 2*1 -----> That is multiplying by 2. See if you can simplify that.

Directrix (directrix):

2 * (1/2) x + 2 * 3y = 2*1 x + 6y = 2

Directrix (directrix):

1/2x + 3y = 1 ---> x + 6y = 2 x + 6y = 2 ---> x + 6y = 2 ------------- Notice that we have the same equation used twice. How do you think that they might intersect. At each point they intersect is a solution.